# How robust are the parameter constraints extending the $\Lambda$CDM model?

### v1

@@ III.1 Base $\Lambda$CDM parameters
Figure 8: Marginalized 1D posteriors for $\Omega_{\mathrm{m}}$, considering different datasets. Colors are the same as in Fig. 1.
Refer to captionRefer to caption
Figure 9: Marginalized 1D posteriors for $\sigma_{8}$, considering different datasets. Colors are the same as in Fig. 1.
Refer to captionRefer to caption
Figure 10: Marginalized 1D posteriors for $S_{8}$, considering different datasets. Colors are the same as in Fig. 1.
Refer to captionRefer to caption
@@
Very similar results to those previously discussed for $\tau_{\rm{reio}}$ are obtained for the amplitude of the primordial power spectrum $A_{\rm s}$, strongly correlated with $\tau_{\rm reio}$: its most accurate measurements arise from Planck, low redshift data does not improve its determination and the largest impact on its error is when the lensing amplitude is a free parameter (see Tab. 10), as $A_{\rm s}$ controls the overall normalization of the CMB power spectrum. The two panels of Figure 6 depict the one-dimensional probability distribution for this parameter for different data sets and fiducial cosmologies.
III.2 Derived parameters
We have discussed so far the main six $\Lambda$CDM parameters. Now we will focus on the derived ones, starting with the Hubble constant, $H_{0}$. Table 4 shows the $95\%$ CL constraints on the Hubble parameter. While the mean value of $H_{0}$ is barely shifted, its errors are increased by one to two orders of magnitude when only CMB data are considered, depending on the data set. This case is completely different from what we have observed so far, i.e. low redshift measurements are crucial for restoring the errors to their default values within the $\Lambda$CDM paradigm, see Tab. 5. Once low redshift probes are included in the analyses, the errors on $H_{0}$ are decreased by more than one order of magnitude. The very large spread in the errors of the Hubble constant is clearly visible from the left panel of Fig. 7. The right panel shows the dilution of such a spread due to BAO observations, which break many of the degeneracies between $H_{0}$ and the cosmological parameters involved in the extended cosmological scenarios considered here. The largest errors and departures on $H_{0}$ are obtained when the dark energy sector is modified. As can be noticed from Tabs. 11 and 12, a much larger value of the Hubble constant is obtained when the dark energy equation of state is a freely varying parameter or a function of redshift for the case of Planck and ACT. In addition, the value of the dark energy equation of state is preferred to be in the phantom region with a significance of $2\sigma$ (in the constant $w$ case). This is a well-known result and indeed it has been proposed as a scenario where to solve the long-standing Hubble constant tension Di Valentino et al. (2016). In the case of SPT, the situation is reversed: a value $w>-1$ is preferred, and therefore the value of the Hubble parameter is considerably smaller than within the minimal $\Lambda$CDM scheme. Nevertheless, also in the SPT case the low redshift data restores mean values and errors of the Hubble constant very close to their $\Lambda$CDM counterparts.
The following derived parameter we discuss here is the matter mass energy density parameter, $\Omega_{\rm m}$. As in the previous case, low redshift measurements also affect in a significant way the errors on the marginalized values, see Tab. 4, and their addition reduces the errors by one order of magnitude in some cases. The very large spread in the mean values and errors of $\Omega_{\rm m}$ can be noticed from the left panel of Fig. 8 and the constraining power of low-$z$ data is visible from the right panel of the very same figure. The largest departure in the mean values of the matter density parameter occurs, as in the previous case, when the dark energy sector freedom is enlarged: the reason for that is simple. As we have seen, for the case of Planck and ACT, the Hubble constant can take very large values when $w<-1$. However, the CMB acoustic peak structure does not allow the quantity $\Omega_{\rm m}h^{2}$ to be very large, and, consequently, the value of $\Omega_{\rm m}$ is required to be considerably smaller than in the canonical $\Lambda$CDM scenario. In the case of SPT the situation is again reversed, and a much larger value of $\Omega_{\rm m}$ is obtained. However, for the three CMB datasets explored here (Planck, ACT and SPT), when low redshift measurements are included, the value of the dark energy equation of state gets much closer to $-1$ and the mean values and errors are almost restored to their default values within the $\Lambda$CDM picture.
Concerning the remaining derived parameters, i.e. the clustering-related ones $\sigma_{8}$ and $S_{8}$, and $r_{\rm drag}$, low redshift measurements play a major role especially in the case of $\sigma_{8}$ and $S_{8}$. As visible from Figs. 9 and 10, the addition of these measurements dilutes the spread in the mean values and errors of these two clustering parameters, ensuring their stability against different possible fiducial cosmologies. Therefore, it is perfectly valuable to use the marginalized values of these derived parameters when testing exotic cosmological scenarios, due to their stability against different possible fiducial cosmologies.
@@ IV Conclusions
Are the $\Lambda$CDM model parameters robust? The answer is yes. After analyzing the model-marginalized limits on the standard $\Lambda$CDM cosmological parameters $\Omega_{\rm c}h^{2}$, $\Omega_{\rm b}h^{2}$, $\theta_{\rm MC}$, $\tau_{\rm reio}$, $n_{s}$ and $A_{s}$ as well as on the derived quantities $H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$ and $r_{\rm drag}$ we notice that even if with CMB data only the mean values and errors have large spreads when computed over a number of possible non-minimal fiducial cosmologies, the addition of low redshift measurements restores their values to their canonical $\Lambda$CDM ones. Our marginalization procedure includes scenarios with massive neutrinos, extra relativistic species, non-minimal dark energy equation of state scenarios, non-zero curvature or a varying lensing amplitude parameter. Our results not only demonstrate the stability of the standard cosmological parameters, but also provide a set of marginalized errors and mean values that can be used as an input for cosmological fitting purposes, and they are more reliable and conservative than those obtained within the minimal $\Lambda$CDM picture.
@@ Acknowledgements.
EDV is supported by a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology). We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. S.G. is supported by the European Union’s Framework Programme for Research and Innovation Horizon 2020 (2014–2020) under Junior Leader Fellowship LCF/BQ/PI23/11970034 by La Caixa Foundation. This work has been supported by the Spanish MCIN/AEI/10.13039/501100011033 grants PID2020-113644GB-I00 and by the European ITN project HIDDeN (H2020-MSCA-ITN-2019/860881-HIDDeN) and SE project ASYMMETRY (HORIZON-MSCA-2021-SE-01/101086085-ASYMMETRY) and well as by the Generalitat Valenciana grants PROMETEO/2019/083 and CIPROM/2022/69. OM acknowledges the financial support from the MCIU with funding from the European Union NextGenerationEU (PRTR-C17.I01) and Generalitat Valenciana (ASFAE/2022/020).
@@ Appendix A Tables
In this Appendix we include the tables with all the results for the cosmological parameters within all the models discussed in our work.

### v2

@@ III.7 Hubble Constant, $H_{0}$
Figure 10: Marginalized 1D posteriors for the matter energy density $\Omega_{\mathrm{m}}$ and the matter clustering parameters $\sigma_{8}$ and $S_{8}$, considering different datasets.Colors are the same as in Fig. 2.
III.7.1 Main correlations with beyond $\Lambda$CDM parameters
When examining the correlation between $H_{0}$ and other cosmological parameters, the most significant increases in the errors and/or deviations from results inferred within a baseline $\Lambda$CDM cosmology occur when modifications are made to the dark energy sector. As can be noticed from Tabs. 10 and 11, a much larger value of the Hubble constant is obtained when the dark energy equation of state is a freely varying parameter or a function of redshift for the case of Planck and ACT. This is because of the well-known geometrical degeneracy among the equation of state, the matter density, and the Hubble constant. As seen from the left bottom panel of Fig. 8, the value of the dark energy equation of state is preferred to be in the phantom region with a significance of $2\sigma$ (in the constant $w$ case). Therefore, changes in the properties or dynamics of dark energy can substantially impact the estimation of $H_{0}$ and its relationship with other parameters. This is a well-known result and indeed it has been proposed as a scenario where to solve the long-standing Hubble constant tension Di Valentino et al. (2016). In the case of SPT, the situation is reversed: a value $w>-1$ is preferred (see the one dimensional probability distribution functions in Fig. 8), and therefore the value of the Hubble parameter is considerably smaller than within the minimal $\Lambda$CDM cosmology. Nevertheless, also in the SPT case the low redshift data restores mean values and errors of the Hubble constant very close to their $\Lambda$CDM counterparts. Overall, as largely expected, the inclusion of low-$z$ datasets becomes crucial when simultaneously varying parameters related to the dark sector and $H_{0}$. To emphasize this aspect, in Fig. 8, we have kept the same $y$-axis scale when plotting the two-dimensional marginalized contours with and without low-$z$ information to highlight the significant reduction in contour size compared to those obtained from CMB-only analyses.
Another significant shift of $H_{0}$ towards smaller values is instead due to the presence of massive neutrinos, that change the angular diameter distance at the last scattering surface and thus introduce a correlation with the Hubble constant. In this case, the shift for all the CMB experiments is in the same direction. Varying the neutrino effective number $N_{\rm eff}$ is again another method to relax the $H_{0}$ constraints, because of their very well known correlation through the effect on the peaks in the damping tail. The one-dimensional posterior distribution function and the two-dimensional correlation between $H_{0}$ and $N_{\rm eff}$ are shown in the top-left panel of Fig. 8 for the different datasets. Also in this case, the final preferred value depends on the CMB observations under consideration: ACT points towards a lower value for $N_{\rm eff}$ and therefore a lower $H_{0}$ also, while SPT has the opposite trend. The very same argument can be used when the spatial curvature of the universe is free to vary. ACT goes in the direction of a closed universe and a lower $H_{0}$ value with error bars increasing up to a factor 5 with respect to the $\Lambda$CDM case, while SPT prefers an open universe and a higher $H_{0}$ value.
Last but not least, other important correlations involve the curvature parameter $\Omega_{k}$ and the lensing amplitude $A_{\rm lens}$, both of which are displayed in Fig. 8 for the different datasets.
III.7.2 Marginalized Constraints
Table 3 shows the $95\%$ CL constraints on the Hubble parameter. While the mean value of $H_{0}$ is barely shifted, its errors are increased by one to two orders of magnitude when only CMB data are considered, depending on the dataset. This case is completely different from what we have observed so far, i.e., low redshift measurements are crucial for restoring the errors to their default values within the $\Lambda$CDM paradigm, see Tab. 4, Fig. 9 and Fig. 8. Once low redshift probes are included in the analyses, the errors on $H_{0}$ are decreased by more than one order of magnitude. The very large spread in the errors of the Hubble constant is clearly visible from the left panel of Fig. 9. The right panel shows the dilution of such a spread due to BAO observations, which break many of the degeneracies between $H_{0}$ and the cosmological parameters involved in the extended cosmological scenarios considered here when only CMB observations are considered.
@@ III.7.2 Marginalized Constraints
III.8 Matter Energy Density $\Omega_{\rm m}$ and Clustering parameters, $\sigma_{8}$ and $S_{8}$
The following derived parameters we discuss are those related to the matter sector of the Universe, namely, the matter energy density, $\Omega_{\rm m}$, and the matter clustering-related parameters $\sigma_{8}$ and $S_{8}$.
III.8.1 Main correlations with beyond $\Lambda$CDM parameters
The most relevant correlations involving the matter density and the clustering parameters are displayed in Fig. 8. In the top right panel, we show the correlation with the effective number of relativistic degrees of freedom. The top left panel highlights correlations with the lensing amplitude, while the bottom left explores the marginalized constraints with the curvature energy density. Finally, the bottom right panel shows the constraints with the dark energy equation of state. Among those, the largest shifts in both the matter density and matter clustering parameters occur when freedom is introduced into the dark energy sector. For datasets such as Planck and ACT, the Hubble constant can take higher values when $w<-1$. However, the structure of the CMB acoustic peaks restricts $\Omega_{\rm m}h^{2}$ from becoming too large. As a result, $\Omega_{\rm m}$ needs to be smaller than in the standard $\Lambda$CDM model. Conversely, in the case of SPT, a much larger value of $\Omega_{\rm m}$ is observed, which is almost unconstrained within the considered prior range. For a similar reason (i.e., its correlation with $H_{0}$), the matter density shifts to higher values when considering massive neutrinos, while the uncertainties relax by a factor of 2-3 compared to the $\Lambda$CDM scenario across all CMB datasets. However, when low-redshift measurements are added for Planck, ACT, and SPT, the value of the dark energy equation of state approaches $w=-1$, bringing the mean values and uncertainties back in line with $\Lambda$CDM expectations. Low-redshift measurements also play a critical role in the constraints on $\sigma_{8}$ and $S_{8}$. Since low-redshift data directly measure these parameters, they help in breaking the degeneracies present in CMB-only analyses. Lastly, varying the neutrino effective number $N_{\rm eff}$ or the spatial curvature $\Omega_{k}$ results in significant shifts: ACT and Planck observations push the parameters towards higher values, while SPT tends to shift them towards lower values.
III.8.2 Marginalized Constraints
The marginalized constraints on the matter-related parameters are summarized in Tab. 3. The posterior distribution functions for $\Omega_{\rm m}$, $\sigma_{8}$, and $S_{8}$, as well as the shifts in the posteriors derived from different models and datasets, are shown in Fig. 10. In the left panel of Fig. 10, we see that when considering only CMB data, there is significant scatter in the mean values and uncertainties of $\Omega_{\rm m}$, $\sigma_{8}$, and $S_{8}$. This is due to the challenge CMB data alone faces in constraining these quantities in extended cosmological models, largely because of correlations with other parameters, as discussed in the previous subsection. However, low-redshift measurements significantly reduce the uncertainties in the marginalized values. In some cases, they reduce the errors by an order of magnitude. The constraining power of low-$z$ data is evident in the right panel of the same figure, where the addition of these measurements narrows the scatter in the mean values and uncertainties of both the matter-energy density and the two clustering parameters, ensuring their stability against different possible fiducial cosmologies. As a result, the marginalized values of these derived parameters, especially when low-redshift data are included, offer a robust basis for testing exotic extensions of the $\Lambda$CDM scenario, due to their stability across various cosmological models.
III.9 The sound horizon, $r_{\rm drag}$
The last parameter we discuss is the value of the sound horizon at the drag epoch, $r_{\rm drag}$. [^5] It is important to note that the sound horizon evaluated at the baryon drag epoch, referred to here as $r_{\rm drag}$, differs from the sound horizon evaluated at recombination because the two epochs are separated by $\Delta z\sim 30$. Since the results in the literature are commonly expressed in terms of $r_{\rm drag}$, we follow this convention.
III.9.1 Main correlations with beyond $\Lambda$CDM parameters
By definition, the sound horizon contains information about the expansion history of the Universe from soon after inflation to before recombination. Therefore, when studying the correlation between the sound horizon and other cosmological parameters, it is no surprise that this parameter is primarily sensitive to quantities that influence the expansion history of the Universe before recombination. The main correlation we pointed out is with the effective number of relativistic degrees of freedom, $N_{\rm eff}$. Specifically, larger values of $N_{\rm eff}>3$ will increase the radiation energy density of the Universe before recombination, leading to a faster expansion rate, $H(z)$. This, in turn, results in a smaller value for the sound horizon. Conversely, $N_{\rm eff}<3$ has the opposite effect, leading to larger values of $r_{\rm drag}$. Due to its critical role in potentially addressing the long-standing Hubble tension, the correlation between the sound horizon and other cosmological parameters has been widely discussed in the literature. We have also highlighted these correlations ourselves in relation to other key cosmological parameters, such as $\theta_{\rm MC}$. Therefore, to avoid repetition of well-established information, we refer to previous subsections and the extensive literature dedicated to this topic for further details.
III.9.2 Marginalized Constraints
The marginalized constraints on the sound horizon at the drag epoch are summarized in Table 3 and visualized in Fig. 11. Several important observations are warranted here. Firstly, we note that the constraints achievable with Planck CMB observations are more than three times tighter compared to those based on ACT or SPT data. Including low-redshift information significantly improves the precision attainable with ACT and SPT, while it has a less pronounced impact on the Planck results. Consequently, when combining CMB data with local probes, the difference in constraining power between different datasets is reduced to a factor of 2. The improvement in the constraining power from local Universe probes is evident when comparing the left and right panels of Fig. 11. This can be explained by the sensitivity of the sound horizon to beyond-$\Lambda$CDM extensions. Despite this improvement, we observe that the marginalized uncertainties remain relatively large for experiments beyond Planck, and even within Planck they are broader than those inferred assuming a $\Lambda$CDM cosmology. This highlights how sensitive constraints on the sound horizon can be to changes in fiducial cosmology. This sensitivity includes not only parameters that affect early-time cosmology (such as $N_{\rm eff}$, discussed in the previous subsection) but also parameters influencing late-time expansion history, particularly for experiments other than Planck. This is evident from Fig. 11, where we observe significant scatter in the one-dimensional probability distribution functions obtained across different fiducial cosmologies.
We conclude with a final remark on the applicability of our marginalized constraints on the sound horizon. The value of $r_{\rm drag}$ inferred by Planck, assuming a baseline $\Lambda$CDM cosmology, is often adopted as a prior in studies that aim to compress the CMB likelihood when testing extended cosmological models (particularly when no changes are expected in the expansion history of the Universe before recombination). However, as highlighted by our analysis, mild deviations from this value are observed even when studying late-time expanded cosmologies due to correlations among parameters. Therefore, we argue that our marginalized constraints offer a safer and more robust choice for distance prior likelihoods. Our constraints are less model-dependent (although still based on extensions of the $\Lambda$CDM model) and more conservative, with larger uncertainties that better capture potential variations across different fiducial cosmologies. This makes them more suitable for extended model tests, avoiding the overly restrictive assumptions of a $\Lambda$CDM framework.
@@ IV Conclusions
In this work, we provide a comprehensive overview of the constraints on cosmological parameters derived from current CMB experiments, both individually and in combination with local Universe probes. We discuss the various correlations observed across different extended cosmological models on a parameter-by-parameter basis, highlighting key correlations and degeneracy patterns between standard and non-standard cosmological parameters and pointing out the most important differences among different datasets and CMB experiments. Ultimately, we present the marginalized posterior distribution functions, where information from all the considered extensions of the $\Lambda$CDM model converges, leading to constraints on the main cosmological parameters that are more robust than those derived from specific cosmological models.
@@
Our study is motivated by several interesting reasons. First and foremost, we aim to address the following question: Are the $\Lambda$CDM model parameter values robust? The answer is yes. After analyzing the model-marginalized limits, obtained considering several extensions of the simplest $\Lambda$CDM scenario, on the standard $\Lambda$CDM cosmological parameters $\Omega_{\rm c}h^{2}$, $\Omega_{\rm b}h^{2}$, $\theta_{\rm MC}$, $\tau_{\rm reio}$, $n_{s}$, and $A_{s}$, as well as on the derived quantities $H_{0}$, $\Omega_{\rm m}$, $\sigma_{8}$, $S_{8}$, and $r_{\rm drag}$, we find that, although the mean values and errors from CMB data alone have large spreads when computed over a number of possible non-minimal fiducial cosmologies, the addition of low-redshift measurements restores their values to the canonical $\Lambda$CDM ones.
Secondly, our marginalization procedure includes extensions of the minimal $\Lambda$CDM model with massive neutrinos, extra relativistic species, non-minimal dark energy equation of state scenarios, non-zero curvature, or a varying lensing amplitude parameter. Our results not only demonstrate the stability of the standard cosmological parameters within the minimal $\Lambda$CDM model and its extensions, but also provide a set of marginalized errors and mean values that are more reliable and conservative than those obtained within the minimal $\Lambda$CDM picture, serving as valuable tools for testing some extended cosmological scenarios.
In particular, we highlight the following findings and their potential applications. Our marginalized constraints on the baryon energy density parameter, $\Omega_{\rm b}h^{2}$, provide a more robust alternative to the traditional value inferred within $\Lambda$CDM cosmology, owing to their reasonably reduced model dependence. In studies focused on late-time cosmology, a prior on $\Omega_{\rm b}h^{2}$ is often adopted based on results from Planck under the $\Lambda$CDM assumption or, alternatively, from BBN predictions, which typically increase uncertainties compared to Planck-based constraints. Our results offer a reasonable alternative to both the BBN prior and the Planck $\Lambda$CDM-based result. Indeed, our marginalized constraint considers a broad range of models and possible effects and, although not entirely model-independent, can certainly be regarded as more conservative estimates than those derived under the $\Lambda$CDM assumption. They are also at least competitive with the BBN prior in terms of model dependence, making them effective for these analyses.
The Planck measurement of the angular scale of the acoustic peaks, $\theta_{\rm MC}$, is frequently used as a distance prior due to its relatively model-independent nature, being closely related to the position of the first acoustic peak in the CMB angular power spectra. $\theta_{\rm MC}$ can effectively serve as an additional BAO point at recombination (i.e., at $z\sim 1090$) when testing cosmological models and compressing CMB likelihoods. However, state-of-the-art analyses typically rely on the value inferred by Planck assuming a $\Lambda$CDM cosmology. Although this parameter is determined with a high level of precision and is not very sensitive to changes in cosmological models, minor deviations can nonetheless arise when considering various beyond-$\Lambda$CDM parameters. In contrast, our marginalized results account for potential effects arising from a variety of fiducial cosmologies, thereby further reducing the model dependence of our estimate. Thus, our marginalized constraints can and should be utilized in analyses, potentially replacing values obtained from single baseline cosmology assumptions, when a full analysis of cosmological data cannot be performed.
Finally, similarly to the previous parameters, the sound horizon constraint derived by Planck, based on the $\Lambda$CDM assumption, is often used as a prior in studies that aim to compress the CMB likelihood when testing extended cosmological models, particularly when no changes in the expansion history of the Universe are expected before recombination. However, our analysis reveals that mild deviations from this value can occur even when considering late-time expanded cosmologies due to parameter correlations. Consequently, we argue that our marginalized constraints represent a safer and more robust choice for distance prior likelihoods. These constraints are less model-dependent and more conservative, with larger uncertainties that better capture potential variations across cosmological models, thus avoiding the restrictive assumptions inherent in the $\Lambda$CDM framework.
@@ Acknowledgements.
S.G. is supported by the European Union’s Framework Programme for Research and Innovation Horizon 2020 (2014–2020) under Junior Leader Fellowship LCF/BQ/PI23/11970034 by La Caixa Foundation and by the Research grant TAsP (Theoretical Astroparticle Physics) funded by Istituto Nazionale di Fisica Nucleare (INFN). This work has been supported by the Spanish MCIN/AEI/10.13039/501100011033 grants PID2020-113644GB-I00 and by the European ITN project HIDDeN (H2020-MSCA-ITN-2019/860881-HIDDeN) and SE project ASYMMETRY (HORIZON-MSCA-2021-SE-01/101086085-ASYMMETRY) and well as by the Generalitat Valenciana grants PROMETEO/2019/083 and CIPROM/2022/69. OM acknowledges the financial support from the MCIU with funding from the European Union NextGenerationEU (PRTR-C17.I01) and Generalitat Valenciana (ASFAE/2022/020). EDV is supported by a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology). We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.